A computer that was purchased for $6000 depreciates at a rate of 13.4% per year. An equation modeling this is, f(x)=6000(0.866)^x, where x is the number of years.
What is the computer worth after 2 years?
When will the value of the computer be $2530?

Answer by Edwin McCravy(20067)

I think your f(x)=6000(0.866)x should be f(x)=6000(0.866)x.

So post it again as f(x)=6000(0.866)^x where the ^ points up where
the exponent is supposed to go.  If you write x down on the line with
everything else it will look like multiplication by x instead of to
the x power.

Edwin